The value of ∫x2(1-x)3/2dx

Grade:12

1 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Suhas,

 

Make the sub 1-x = t^2, so that dx = -2tdt.

And the integral becomes ∫[(1-t^2)^2](t^3)[-2tdt].

Open the brackets, and you will get a directly integrable function.

 

Hope it helps.

 

Best Regards,

Ashwin (IIT MadraS).